The management of Brinkley Corporation is interested in using simulation to estimate the profit per unit for a new product. The selling price for the product will be $45 per unit. Probability distributions for the purchase cost, the labor cost, and the transportation cost are estimated as follows: Procurement Cost ($) Probability Labor Cost ($) Probability Transportation Cost ($) Probability 10 0.25 20 0.10 3 0.75 11 0.45 22 0.25 5 0.25 12 0.30 24 0.35 25 0.30 a) Construct a simulation model to estimate the mean profit per unit. b) Management believes that the project may not be sustainable if the profit per unit is less than $5. Use simulation to estimate the probability that the profit per unit will be less than $5....Need clear explaination with formulas. please urgent
Answer #1

ANSWER

(A) : Simulation model

Procurement cost($) Probability Labour cost ($) Probability Transportation cost($) Probability
10 0.00 but less than 0.25 20 0.00 but less than 0.10 3 0.00 but less than 0.75
11 0.25 but less than 0.70 22 0.10 but less than 0.35 5 0.75 but less than 1.00
12 0.70 but less than 1.00 24 0.35 but less than 0.70
25 0.70 but less than 1.00

based on simulation model mean profit = 45 - 11 - 24 - 5 = $ 5/ unit and 45 - 10 - 25 - 3 = $ 7 / unit

(B) : As calculated above, based on simulation model, minimum profit is $ 5/ unit. Hence management's belief of non-sustainability of project is right. Less than $ 5 / unit profit scenario is unacceptably low.

-------------------------------------

DEAR STUDENT,

IF YOU HAVE ANY QUERY ASK ME IN THE COMMENT BOX,I AM HERE TO HELPS YOU.PLEASE GIVE ME POSITIVE RATINGS

*****************THANK YOU***************